# SliM-LLM: Salience-Driven Mixed-Precision Quantization for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) achieve remarkable performance in natural language understanding but require substantial computation resources and memory footprint. Post-training quantization (PTQ) is a powerful compression technique extensively investigated for its effectiveness in reducing memory usage and improving the inference efficiency of LLMs. However, existing PTQ methods are still not ideal in terms of accuracy and efficiency, especially with below 4 bit-widths. Standard PTQ methods using group-wise quantization suffer difficulties in quantizing LLMs accurately to such low-bit, but advanced methods remaining high-precision weights element-wisely are hard to realize its theoretical hardware efficiency. This paper presents a Salience-Driven Mixed-Precision Quantization scheme for LLMs, namely SliM-LLM. The scheme exploits the salience distribution of LLM weights to determine optimal bit-width and quantizers for accurate LLM quantization, while aligning bit-width partition to quantization groups for compact memory usage and fast integer computation on hardware inference. Specifically, the proposed SliM-LLM mainly relies on two novel techniques: (1) Salience-Determined Bit Allocation utilizes the clustering characteristics of salience distribution to allocate the bit-widths of each quantization group. This increases the accuracy of quantized LLMs and maintains the inference efficiency high; (2) Salience-Weighted Quantizer Calibration optimizes the parameters of the quantizer by considering the element-wise salience within the group. This balances the maintenance of salient information and minimization of errors. Comprehensive experiments show that SliM-LLM significantly improves the accuracy of various LLMs at ultra-low 2-3 bits, e.g., 2-bit LLaMA-7B achieves a 5.5-times memory-saving compared to the original model on NVIDIA A800 GPUs, and 48% decrease of perplexity compared to the state-of-the-art gradient-free PTQ method. Moreover, SliM-LLM+, which is integrated from the extension of SliM-LLM with gradient-based quantizers, further reduces perplexity by 35.1%. We highlight that the structurally quantized features of SliM-LLM exhibit remarkable versatility and promote improvements in the accuracy of quantized LLMs while keeping inference efficiency on hardware. Our code is available at https://github.com/Aaronhuang-778/SliM-LLM.
@@ 1 Introduction
Large language models (LLMs) have exhibited exceptional performance across a wide array of natural language benchmarks [3, 48, 19, 2]. Notably, LLaMA [41] and GPT [3] series have significantly contributed to the ongoing evolution of LLMs towards universal language intelligence. The powerful language understanding capabilities of LLMs have been transferred to multi-modal domains [25, 1, 40, 50], laying the foundation for artificial general intelligence [4]. Despite these significant achievements, the substantial computational and memory requirements of LLMs pose efficiency challenges for real-world applications and deployments, particularly in resource-constrained environments. For example, the latest LLaMA-3-70B111 https://github.com/meta-llama/llama3 model, with its 70 billion parameters, requires over 150GB of storage and a minimum of two NVIDIA A800 GPUs, each with 80GB of memory, for inference [21].
To reduce the computation burden, post-training quantization (PTQ), as an efficient and effective compression approach [11], has also been explored and proven successful in quantizing the weights of pre-trained LLMs [16, 20, 26, 38, 23, 6]. Faced with the dilemma of scaled-up LLMs and the limited computation resources, there is an urgent need for more aggressive compression [20, 43]. However, despite considerable efforts, significant performance degradation still occurs in low bit-width scenarios ($\leqslant$ 3-bit). To maintain the performance, unstructured mixed-precision quantization schemes [37, 20, 12] or specialized transformation computations [6, 43, 13, 7] are necessary. Yet, these approaches impose additional burdens on hardware during the inference of quantized LLMs, suffering memory overhead of element-wise bit-maps and computation overhead of codebook decoding and bit-map addressing (even preventing efficient integer computation). Moreover, even though fine-tuning can improve the accuracy of quantized LLMs, it increases the overfitting risk and requires expensive computation resources and a long time [26, 5]. Consequently, ensuring the accuracy of LLMs while maintaining efficiency during deployment remains a significant challenge for current PTQ approaches.
@@
Figure 1: (a) The perplexity ($\downarrow$) of existing low-bit PTQ methods of LLaMA at 2-bit. Solid-line indicates structured quantization methods. (b) Compare PTQ methods with gradient quantizer at 3-bit. (c) Features of current low-bit quantization methods. C denotes codebook-based, S is statistic-based, and G represents gradient-based quantizers.
This paper presents the Salience-Driven Mixed-Precision LLM (SliM-LLM) framework, an accurate and hardware-efficient PTQ method for LLMs ($\leqslant$ 3-bit). SliM-LLM can be seamlessly integrated into existing advanced PTQ pipelines [16, 38], as a plug-and-play approach with mixed-precision computing for improved performance (Fig. 1). Our approach builds on the observation that not all parameters are equally important [12, 20, 37]. Specifically, a subset of salient weights significantly influences an LLM’s capabilities and tends to be concentrated in specific channels. During structured group-wise quantization, the uneven distribution of these salient channels leads to differential importance across various groups. Based on this finding, we design a structured mixed-precision quantization approach for LLMs. First, we develop a novel Salience-Determined Bit Allocation (SBA) method to allocate the optimal bit-width configuration for each structured group based on the salience distribution, minimizing the weight output relative entropy. By implementing bit-width compensation constraints, SBA maintains the average bit-width, while improving the low-bit performance. Next, we introduce the Salience-Weighted Quantizer Calibration (SQC), which amplifies the awareness of locally salient weights, preventing the degradation of sensitive information within groups. SQC works collaboratively with SBA, exploiting the local and global salience of weights to preserve the performance of LLMs after quantization. Notably, SliM-LLM does not rely on fine-tuning processes, efficiently deploying weight quantization on various LLMs. Moreover, compared to the unstructured mixed-precision methods [37, 12, 20], SliM-LLM incurs no additional bits and computation overhead. We also deploy SliM-LLM on the application-level inference tool 222 https://github.com/AutoGPTQ/AutoGPTQ for LLMs, facilitating mixed-precision inference on graphics processing units (GPUs) with high performance.
@@
Experiments show that for various LLM families, SliM-LLM surpasses existing PTQ methods on diverse benchmarks as a plug-and-play unit, particularly in low-bit scenarios. Using GPTQ as the backbone, SliM-LLM improves the perplexity scores of 2-bit LLaMA-13B and LLaMA2-13B on WikiText2 [30] from 20.44 and 28.14 to 8.87 and 9.41, denoting performance improvements of over 56%, respectively. SliM-LLM even outperforms other unstructured mixed-precision PTQ methods, such as PB-LLM [37], APTQ [18] and LLM-MQ [24], in a deployment-friendly manner, showcasing its superior low-bit accuracy and efficiency. Moreover, we integrate SliM-LLM into OmniQuant [38] and obtain SliM-LLM+ through gradient optimization to further improve quantization quality.
@@ 2 Related Work
Large Language Models (LLMs) have been significantly developed in diverse natural language processing domains, establishing a prominent paradigm in these fields [4, 5, 51, 3, 41]. Nevertheless, the exceptional success of LLMs depends on massive parameters and computations, posing significant challenges for deployment in resource-constrained environments. Consequently, research into the compression of LLMs has emerged as a promising field. Existing compression techniques for LLMs primarily include low-bit quantization, pruning, distillation, and low-rank decomposition [46, 17, 16, 44, 38, 6, 52, 15, 20, 33, 7]. Among these technologies, low-bit quantization gains remarkable attention, for efficiently reducing the model size without change of network structure[52, 51, 5].
Quantization of LLMs can be generally divided into quantization-aware training (QAT) [27] and post-training quantization (PTQ) [44, 16, 38]. QAT, by employing a retraining strategy based on quantized perception, better preserves the performance of quantized models. LLM-QAT [27] addresses the data obstacle issue in QAT through data-free distillation. However, for LLMs with huge size of parameters, the cost of retraining is extremely inefficient[5]. Therefore, PTQ has become a more efficient choice for LLMs. For instance, LLM.int8() [27] and ZeroQuant [47] explore the quantization strategies for LLMs in block-wise, which is a low-cost grouping approach that reduces hardware burden. Smoothquant [44] scales weight and activation to decrease the difficulty of quantization. Subsequently, AWQ [26] and OWQ [23] also propose scaling transformations on outlier channels of weight to preserve their information representation capacity. GPTQ [16] reduces the group quantization error of LLMs through Hessian-based error compensation [14], achieving commendable quantization performance at 3-bit. OmniQuant [38] introduces a learnable scaling quantizer to reduce quantization errors in an output-aware manner. To enhance the accuracy of LLMs at 3-bit, APTQ [18] allocates different bit-width to different transformer blocks based on Hessian-trace, enhancing the accuracy of LLMs 3-bit. To achieve LLM quantization at ultra-low bit-width, recent novel efforts such as QuIP [6], QuIP# [43], and AQLM [13] promote quantization performance at 2-bit through learnable codebooks or additional fine-tuning. Meanwhile, approaches like SpQR[12], PB-LLM [37], and BiLLM [20] employ finer-grained partitioning for grouped quantization with unstructured mixed-precision for weights, further improving the PTQ performance. However, existing low-bit methods still rely on special structures and fine-grained grouping to ensure accuracy, which increases the difficulty of hardware deployment. Additionally, the extra fine-tuning training may pose a risk of domain-specific overfitting and undermine the efficiency of PTQ [26].
@@ 3 SliM-LLM
This section introduces a mixed-precision quantization technique called SliM-LLM, designed to overcome the performance and inference efficiency bottlenecks in mixed-precision frameworks. To address these challenges, we devise two novel strategies for LLMs, including the use of Salience-Determined Bit Allocation (SBA) based on global salience distribution to determine group bit-widths, and Salience-Weighted Quantizer Calibration (SQC) to enhance the perception of locally important weight information. We introduce SBA and SQC in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.1 Preliminaries
Quantization Framework. We first present the general uniform quantization process of LLMs according to common practice [27, 38, 1]. The quantization process requires mapping float-point weights distributed within the interval $[w_{\mathrm{min}},w_{\mathrm{max}}]$ to an integer range of $2^{N}$, where $N$ is the target bit-width. The quantization function for weight $\boldsymbol{w}_{f}\in\mathbb{R}^{n\times m}$ follows:
$\vspace{-0.8pt}\hat{\boldsymbol{w}}_{q}=\operatorname{clamp}(\lfloor\frac{\boldsymbol{w}_{f}}{\Delta}\rceil+z,0,2^{N}-1),~{}\Delta=\frac{w_{\mathrm{max}}-w_{\mathrm{min}}}{2^{N}-1},~{}z=-\lfloor\frac{w_{\mathrm{min}}}{\Delta}\rceil\vspace{-0.8pt}$ (1)
@@
$\hat{\boldsymbol{w}}_{b}=\operatorname{sign}(\boldsymbol{w}_{f}),~{}\operatorname{sign}(w)=\begin{cases}1&\text{if $w\geq 0$},\\ -1&\text{others}.\end{cases},~{}\alpha=\frac{1}{l}||\boldsymbol{w}_{f}||_{\ell 1}$ (2)
where $\hat{\boldsymbol{w}}_{b}$ is binary result. $\alpha$ denots binarization scales and $l$ is the number of elements in weight [34]. We can formalize the per-layer loss in PTQ, following the common practice [31, 16]:
$\mathcal{L}(\hat{\boldsymbol{w}}_{f})=||\boldsymbol{x}\boldsymbol{w}_{f}^{\top}-\boldsymbol{x}\hat{\boldsymbol{w}}_{f}^{\top}||^{2}=\operatorname{tr}((\hat{\boldsymbol{w}}_{f}-\boldsymbol{w})\boldsymbol{H}(\hat{\boldsymbol{w}}_{f}-\boldsymbol{w})^{\top})$ (3)
where $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ denotes the input vectors from calibration dataset, $\hat{\boldsymbol{w}}_{f}\in\mathbb{R}^{n\times m}$ is dequantized weight from quantization result in Eq. (1) or Eq. (2), and $\boldsymbol{H}=\frac{1}{P}\sum_{k=1}^{P}\boldsymbol{x}^{[k]}\boldsymbol{x}^{[k]\top}$ is proxy Hessian matrix by Levenberg-Marquardt approximation [29, 14] from a set of input activations.
Parameter Salience. In LLMs, the importance of each individual element in weight matrix is various [12, 15]. According to Eq. (3), the impact of quantizing a single element on the model’s output loss differs. Elements that significantly influence the loss are termed salient weights. Consequently, we follow the SparseGPT [15] to define the salience of each element as:
Definition 1.
In the quadratic approximation of the loss as expressed in Eq. (3), we give the Hessian matrix $H\in\mathbb{R}^{m\times m}$ generated by $\frac{1}{P}\sum_{k=1}^{P}\boldsymbol{x}^{[k]}\boldsymbol{x}^{[k]^{\top}}$ for a weight matrix, the removal of the element at $(i,j)$ induces an error $\delta_{i,j}=\frac{w_{i,j}^{2}}{[\boldsymbol{H}^{-1}]_{j,j}^{2}}$ to the output matrix for linear projection in LLMs.
where $[\boldsymbol{H}^{-1}]_{jj}$ denotes the $j^{th}$ diagonal entry for the inverse Hessian, and $\boldsymbol{H}^{-1}$ can be efficiently calculated through Cholesky decomposition [22]. According to Definition. 1, we map the elimination error $\delta_{ij}$ to the salience measure of each weight element in LLMs, representing the impact of different weights on the output loss and the language capabilities, which also leads the generation of mixed-precision quantization strategies [12, 37, 20, 24] for LLMs. However, existing mixed-precision solutions require the discrete allocation of bit-widths across the entire weight matrix, which imposes a significant burden on hardware computations, thereby affecting the inference efficiency.

### v2

@@ Abstract
Post-training quantization (PTQ) is an effective technique for compressing large language models (LLMs). However, while uniform-precision quantization is computationally efficient, it often compromises model performance. To address this, we propose SliM-LLM, a salience-driven mixed-precision quantization framework that allocates bit-widths at the group-wise. Our approach leverages the observation that important weights follow a structured distribution and introduces two key components: 1) Salience-Determined Bit Allocation adaptively assigns bit-widths to groups within each layer based on their salience; and 2) Salience-Weighted Quantizer Calibration optimizes quantizer parameters by incorporating element-level salience. With its structured partitioning, SliM-LLM provides a hardware-friendly solution that matches the efficiency of uniform quantization methods while improving accuracy. Experiments show that SliM-LLM achieves superior performance across various LLMs at low bit-widths. For example, a 2-bit quantized LLaMA-7B model reduces memory usage by nearly 6x compared to the floating-point baseline, decreases perplexity by 48% compared to state-of-the-art gradient-free PTQ methods, and maintains GPU inference speed. Additionally, the extended version, SliM-LLM+, which incorporates gradient-based quantization, further reduces perplexity by 35.1%. Our code is available at https://github.com/Aaronhuang-778/SliM-LLM.
Machine Learning, ICML
@@ 1 Introduction
LLMs have demonstrated remarkable performance on various natural language benchmarks (Brown et al., 2020; Hendrycks et al., 2020). Models like LLaMA (Touvron et al., 2023a) and GPT (Brown et al., 2020) have driven progress toward universal language intelligence. Their capabilities have also extended to multi-modal domains (Li et al., 2024b; Achiam et al., 2023; Team et al., 2023; Zhang et al., 2023; Huang et al., 2024b), advancing efforts toward artificial general intelligence (AGI) (Bubeck et al., 2023). However, their high computational and memory demands remain a significant challenge for practical deployment.
@@
Figure 1: (a) The perplexity ($\downarrow$) of existing low-bit PTQ methods of LLaMA at 2-bit. Solid-line indicates methods with structured quantization group. (b) Compare PTQ methods with gradient quantizer at 3-bit. (c) Features of current low-bit quantization methods. C denotes codebook-based, S is statistic-based, and G represents gradient-based quantizers.
To address resource constraints of LLMs, PTQ has emerged as an efficient yet effective compression technique (Dettmers et al., 2022), showing success in quantizing the weights of pre-trained LLMs (Frantar et al., 2022; Lin et al., 2023; Shao et al., 2023; Lee et al., 2023; Chee et al., 2024). As LLMs continue to scale, the demand for more aggressive low-bit compression becomes critical due to limited computational and storage resources in application (Huang et al., 2024a; Tseng et al., 2024). However, significant performance degradation remains a challenge in low-bit scenarios ($\leqslant$ 3-bit). To mitigate this, unstructured mixed-precision quantization (Shang et al., 2023; Huang et al., 2024a; Dettmers et al., 2023) and vector quantization (Chee et al., 2024; Tseng et al., 2024; Egiazarian et al., 2024) methods have been developed to preserve performance. While these approaches have advanced the field, they are often hardware-unfriendly, introducing extra storage requirements such as storing bitmaps or code indices. This creates a bottleneck, limiting further reductions in memory and computational demands during deployments.
This paper presents a Salience-Driven Mixed-Group LLM (SliM-LLM) framework, an accurate and inference-efficient PTQ method for LLMs ($\leqslant$ 3-bit). Our approach is grounded in the key observation that salient or important weights, which are critical to model performance, exhibit a structured distribution, often clustering within certain channels (see Sec. 3.2.1 and Fig. 3). This insight, largely overlooked by prior research (Frantar et al., 2022), forms the basis for designing SliM-LLM as a structured, hardware-friendly mixed-precision low-bit method. It preserves performance through two key designs that retain important weights at both the global group and local element levels. First, we develop a novel Salience-Determined Bit Allocation (SBA) method, which adaptively assigns bit-widths to each quantization group based on their group-level salience ranking. The allocation strategy is optimized to reduce activation reconstruction errors. By applying higher precision to more important groups and reducing the bit-widths for less critical ones, SBA achieves a low average bit-widths while enhancing the overall performance of LLMs. Next, we introduce the Salience-Weighted Quantizer Calibration (SQC), which enhances sensitivity to locally salient weights, ensuring that critical information within groups is preserved. SQC works collaboratively with SBA, exploiting the local and global salience of weights to preserve the performance of LLMs after quantization. Unlike element-wise mixed-precision methods (Shang et al., 2023; Dettmers et al., 2023; Huang et al., 2024a), SliM-LLM is inherently structured, eliminating additional bit or computational overhead while preserving high performance. This is further demonstrated through our deployment of SliM-LLM in an application-level inference tool 111 https://github.com/AutoGPTQ/AutoGPTQ for LLMs, enabling efficient mixed-precision inference on GPUs with consistently strong performance.
@@
Experiments show that for various LLM families, SliM-LLM surpasses existing training-free PTQ methods on diverse benchmarks, particularly in low-bit scenarios. Using GPTQ as the backbone, SliM-LLM improves the perplexity scores of 2-bit LLaMA-13B and LLaMA2-13B on WikiText2 (Merity et al., 2016) from 20.44 and 28.14 to 8.87 and 9.41, denoting performance improvements of over 56%, respectively. SliM-LLM even outperforms other element-wise mixed-precision PTQ methods, such as PB-LLM (Shang et al., 2023), APTQ (Guan et al., 2024) and LLM-MQ (Li et al., 2024a), in a deployment-friendly manner, showcasing its superior low-bit accuracy and efficiency. We also integrate SliM-LLM into OmniQuant (Shao et al., 2023) and obtain SliM-LLM+ through gradient optimization to further improve quantization quality. Moreover, the group-wise mixed-precision strategy can smoothly be adapted to existing quantization-aware training (QAT) (Liu et al., 2023), fine-tuning based (Guo et al., 2023; Liao & Monz, 2024; Dettmers et al., 2024), or codebook-based (Chee et al., 2024; Egiazarian et al., 2024; Tseng et al., 2024) LLMs compression methodologies. This structure of weight salience introduces a new practical view of compression for LLMs.
@@ 2 Related Work
Large Language Models have been significantly developed in diverse natural language processing domains, establishing a prominent paradigm in these fields (Bubeck et al., 2023; Chang et al., 2024; Zhao et al., 2023; Brown et al., 2020; Touvron et al., 2023a). Nevertheless, the exceptional success of LLMs depends on massive parameters and computations, posing significant challenges for deployment in resource-constrained environments. Consequently, research into the compression of LLMs has emerged as a promising field. Existing compression techniques for LLMs primarily include quantization, pruning and distillation (Xu et al., 2023; Ganesh et al., 2021; Frantar et al., 2022; Xiao et al., 2023a; Shao et al., 2023; Chee et al., 2024; Zhu et al., 2023; Frantar & Alistarh, 2023; Huang et al., 2024a; Qin et al., 2024). Low-bit quantization has received notable attention for efficiently reducing the size of the model without changing the structure of the network(Zhu et al., 2023; Zhao et al., 2023; Chang et al., 2024).
Quantization of LLMs can generally be categorized into Quantization-Aware Training (QAT) and PTQ. PTQ has emerged as a more practical alternative. Techniques such as LLM.int8()(Liu et al., 2023) and ZeroQuant(Yao et al., 2022) introduce block-wise quantization, a cost-effective grouping method that reduces hardware overhead. Further advancements, including AWQ (Lin et al., 2023) and OWQ (Lee et al., 2023), apply scaling transformations to outlier weight channels, preserving their representational capacity. GPTQ (Frantar et al., 2022) minimizes group quantization errors using Hessian-based error compensation (Frantar & Alistarh, 2022), achieving notable performance at 3-bit quantization. OmniQuant (Shao et al., 2023) introduces a learnable scaling quantizer to mitigate quantization errors in an output-aware manner. For ultra-low bit-widths quantization, approaches such as QuIP (Chee et al., 2024), QuIP#(Tseng et al., 2024), and AQLM(Egiazarian et al., 2024) improve 2-bit quantization performance through matrix decomposition with learnable codebooks and fine-tuning. Recent works (Qin et al., 2024; Liao & Monz, 2024; Dettmers et al., 2024; Guo et al., 2023) have further refined quantization techniques by leveraging parameter-efficient fine-tuning (PEFT), enabling enhanced performance through additional parameter learning.
Mixed-Precision Quantization leverages the varying importance and redundancy of model parameters by assigning different bit-widths to each component. In traditional visual networks, HAWQ V2 (Dong et al., 2020) and HAWQ V3 (Yao et al., 2021) optimize bit-widths allocation on a layer-wise basis using Hessian analysis and Integer Linear Programming (ILP). Similarly, OMPQ (Ma et al., 2023) employs network orthogonality instead of Hessian for bit-widths optimization. For large language models (LLMs), APTQ (Guan et al., 2024) extends HAWQ’s approach by allocating mixed bit-widths to transformer blocks based on Hessian trace, achieving improved accuracy for 3-bit quantization. However, block-wise or layer-wise mixed-precision strategies at 2-bit fail to maintain performance after compression. To address this, recent methods such as SpQR (Dettmers et al., 2023), PB-LLM (Shang et al., 2023), and LLM-MQ (Li et al., 2023) adopt finer-grained grouping with element-wise mixed-precision for more accurate weight quantization. Despite their improvements, these low-bit approaches rely heavily on fine-grained partitioning, which imposes significant challenges for real hardware deployment and inference speed.
@@ 3 SliM-LLM
This section introduces a structured mixed-precision quantization method, SliM-LLM, to overcome the accuracy and efficiency bottlenecks of mixed-precision LLMs. We devise two novel strategies, including the use of Salience-Determined Bit Allocation (SBA) based on global salience distribution to determine group bit-widths, and Salience-Weighted Quantizer Calibration (SQC) to enhance the perception of locally important weight information. We introduce SBA and SQC in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.1 Preliminaries
Quantization Framework. We first present the general uniform quantization process of LLMs according to common practice (Liu et al., 2023; Shao et al., 2023; Achiam et al., 2023). The quantization process requires mapping float-point weights distributed within the interval $[w_{\mathrm{min}},w_{\mathrm{max}}]$ to an integer range of $2^{N}$, where $N$ is the target bit-widths. The quantization function for weight $\boldsymbol{w}_{f}\in\mathbb{R}^{n\times m}$ follows:
$\vspace{-0.1in}\left\{\begin{array}[]{lr}\hat{\boldsymbol{w}}_{q}=\operatorname{clamp}(\lfloor\dfrac{\boldsymbol{w}_{f}}{s}\rceil+z,0,2^{N}-1),\\ s=\dfrac{w_{\mathrm{max}}-w_{\mathrm{min}}}{2^{N}-1},z=-\lfloor\dfrac{w_{\mathrm{min}}}{s}\rceil\end{array}\right.\vspace{-0.8pt}$ (1)
@@
$\begin{array}[]{lr}\hat{\boldsymbol{w}}_{b}=\operatorname{sign}(\boldsymbol{w}_{f}),\alpha=\frac{1}{l}||\boldsymbol{w}_{f}||_{\ell 1}\\ \operatorname{sign}(w)=\begin{cases}1&\text{if $w\geq 0$},\\ -1&\text{others}.\end{cases}\end{array}$ (2)
where $\hat{\boldsymbol{w}}_{b}$ is binary result. $\alpha$ denots binarization scales and $l$ is the number of elements in weight (Qin et al., 2023), used for dequantization through $\alpha\hat{\boldsymbol{w}}_{b}$. We can formalize the per-layer loss in PTQ, following the common practice (Nagel et al., 2020; Frantar et al., 2022):
$\mathcal{L}(\hat{\boldsymbol{w}}_{f})=||\boldsymbol{x}\boldsymbol{w}_{f}^{\top}-\boldsymbol{x}\hat{\boldsymbol{w}}_{f}^{\top}||^{2}\approx\operatorname{tr}((\hat{\boldsymbol{w}}_{f}-\boldsymbol{w})\boldsymbol{H}(\hat{\boldsymbol{w}}_{f}-\boldsymbol{w})^{\top})$ (3)
where $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ denotes the input vectors from calibration dataset, $\hat{\boldsymbol{w}}_{f}\in\mathbb{R}^{n\times m}$ is dequantized weight from quantization result in Eq. (1) or Eq. (2), and $\boldsymbol{H}=\frac{1}{P}\sum_{k=1}^{P}\boldsymbol{x}^{[k]^{\top}}\boldsymbol{x}^{[k]}$ is proxy Hessian matrix by Levenberg-Marquardt approximation (Marquardt, 1963; Frantar & Alistarh, 2022) from a set of input activations.
Parameter Salience. In LLMs, the importance of each element in the weight matrix is various (Dettmers et al., 2023; Frantar & Alistarh, 2023). According to Eq. (3), quantizing different elements causes different impacts on the model’s output loss. Elements that significantly influence the loss are termed salient weights. Consequently, we follow the SparseGPT (Frantar & Alistarh, 2023) to define the salience of each element as:
Definition 3.1.
In the quadratic approximation of the loss as expressed in Eq. (3), we give the Hessian matrix $H\in\mathbb{R}^{m\times m}$ generated by $\frac{1}{P}\sum_{k=1}^{P}\boldsymbol{x}^{[k]^{\top}}\boldsymbol{x}^{[k]}$ for a weight matrix, the removal of the element at $(i,j)$ induces an error $\delta_{i,j}=\frac{w_{i,j}^{2}}{[\boldsymbol{H}^{-1}]_{j,j}^{2}}$ to the output matrix for linear projection in LLMs.
where $[\boldsymbol{H}^{-1}]_{jj}$ denotes the $j^{th}$ diagonal entry for the inverse Hessian, and $\boldsymbol{H}^{-1}$ can be efficiently calculated through Cholesky decomposition (Krishnamoorthy & Menon, 2013). According to Definition. 3.1, we map the elimination error $\delta_{ij}$ to the salience measure of each weight element in LLMs, representing the impact of different weights on the output loss and the language capabilities, which also leads the generation of mixed-precision quantization strategies (Dettmers et al., 2023; Shang et al., 2023; Huang et al., 2024a; Li et al., 2024a) for LLMs. However, existing mixed-precision solutions require the discrete allocation of bit-widths across the entire weight matrix, which imposes a significant burden on hardware computations, thereby affecting the inference efficiency.